justjm:

I'm supposed to write the following as a definite integral and evaluate it, can someone check it for me?

justjm:

\( \lim_{x \rightarrow \infty} \displaystyle\sum\limits_{k=1}^n (\sqrt{(1+\frac{3k}{n})}\cdot\frac{3}{n}) \) My answer: \(\int_{1}^{4}\sqrt{x}dx = \frac{14}{3}\)

Nnesha:

looks good.

Nnesha:

"n --> infinity" typo..

justjm:

Thanks for the confirmation (: